Fermi National Laboratory
 
Operation Summary for January 2, 2012 - January 4, 2012
 

Monday January 2

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, a MI RF station (MIRF12) tripped off for the first of four times in 22 minutes; it reset. Operations established beam to Meson FTBF experiment T-1015 at 10 AM. MIRF12 tripped off at 3:21 PM and would not reset remotely; Operators reset it locally.

During the evening shift, at 4:14 PM, the MiniBooNE beam permit tripped off for the first of three times in a little over an hour; it reset. At 5:26 PM, BRF19 tripped off or the first of four times in 34 minutes; it reset. Operations disabled T-1019 beam at 10:08 PM.

Tuesday January 3

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 1:49 AM, the Booster West Anode power supply tripped off for the first of two times in nine minutes; it reset. At 2:50 AM, BRF17 tripped off for the first of five times in two hours and 44 minutes; it reset. Operators loaded protons into the Recycler at 7:59 AM.

During the day shift, Operations established beam to T-1019 at 10:17 AM. Muon Rings personnel began a study at 2:01 PM.

During the evening shift, the Muon Rings study ended at 5:50 PM. Operations disabled beam to T-1019 at 10:01 PM; this ended T-1019’s run. At 10:54 PM, the MiniBooNE beam permit tripped off when a Switch magnet’s power supply (MBEX) drifted low; Operators power cycled the supply. Around 11 PM, the I- Source’s output began to drop off; Operators worked at stabilizing and recovering it.

Wednesday January 4

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 2:05 AM, an I- Source vacuum pump power supply tripped off; it reset. A Muon Rings person began a study at 6:39 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo approved of the plans by Booster and Muon Rings personnel to conduct studies. There will be a four-hour shutdown for MI to investigate and repair (if possible in the time allowed) its vacuum valve problem, on Thursday (1/5/12), starting with beam off at 6:30 AM and with beam returning around 10:30 AM. Other machine maintenance will also be conducted during this time.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/04/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory